Suppose the table below shows the prices and quantities of the output produced by an economy in two years. If Year 1 is the base year, then what is the inflation rate in the GDP deflator?

1
Expert's answer
2021-02-11T13:07:51-0500
"Solution"

To calculate the inflation rate in the GDP Deflator, lets consider the table below,



Taking 2000 as the base year, then this can be obtained as below


"Nominal\ GDP\ 2000 = {( P_{cars}2000 \times Q_{cars}2000 )} +{ ( P_{bread}
2000 \times Q_{bread}
2000 )}
 =11,500\\"


"Nominal\ GDP\ 2010 = { P_{cars}\ 2010 \times\ Q_{cars}2010 } + { P_{bread}\ 
2010\times Q_{bread}\ 2010}=16,000,000"

"Real\ GDP\ 2010 ={ ( P_{cars}2000 * Q_{cars}2010 )} +{ ( P_{bread}\ 2000\times\ Q_{bread}\ 2010)}=10,800,000\\"

"Therefore,\ Price\ Deflator\ 2010 = (\frac{Nominal\ GDP\ 2010}{Real\ GDP\ 2010 })\times\ 100= 
148\%" , which is the inflation rate.